what is the answer to 5+7x=11+9x could you show me through it?
Answer by algebrahouse.com(1659)

5 + 7x = 11 + 9x
7x = 6 + 9x {subtracted 5 from both sides}
-2x = 6 {subtracted 9x from both sides}
x = -3 {divided both sides by -2}
www.algebrahouse.com